Wednesday, April 19, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #

Agenda

  1. Status update from detector owners, calibrations, efficiencies, etc.
  2. Review systems owners and update information
  3. Commissioning plan updates from detector systems
  4. Address system specific points of interest for the Fall 2017 run

Minutes

General

  • Joerg Rheinhold mentioned that the field at the aerogel PMT's should be simulated/measured
    • Steve saw no field with a probe
    • A logbook entry is needed with simulation calculations and measurements
  • Mark showed some slides regarding the rates in the SHMS
    • Currently we have the maximum accidental rate
    • The question is weather or not we need to modify the current trigger logic
    • This is a very invasive procedure

Beam Line

  • No update provided

Trigger/DAQ

  • No update provided

HMS

6 GeV Drift Chambers

  • Carlos showed some slides regarding the HMS drift chamber per wire efficiencies
  • Two planes in DC2 had an incorrect mapping, when fixed the efficiencies look good

12 GeV Drift Chambers

  • The chambers are the same structure as SHMS and they are built, tested, and stable
  • Each chambers has a couple of wires with no signals, the problem is likely internal
    • Per chamber there are 2 wires that are dead, one is in the corner
  • If necessary the chambers could be opened up and the wires diagnosed and fixed
  • Cables need to be made and tested
  • Brad Sawatzky has a DAQ which can test a single chamber and if needed then he could integrate another fast bus crate to test both simultaneously
  • We need 76 discriminator cards for the new chambers we need to see how many are available
    • There are 72 in HMS currently

Hodoscopes

  • Simona showed some slides regarding the first stab at the calibrations
  • More tweaks are required, and more diagnostic plots to come

Aerogel

  • No update provided

Heavy Gas Cherenkov

  • No update provided

Shower

  • Vardan reported that the HMS calorimeter was calibrated utilizing run 303 and is close to previous HMS/SOS analyses

SHMS

GEM

  • The GEM has been installed between the SHMS drift chambers, see the logbook entry:
  • The GEM DAQ has been installed in the SHMS electronics hut
  • Work is ongoing to setup the DAQ so that it can eventually be configured into the Hall C DAQ

Nobel Gas Cherenkov

  • No update provided

Drift Chambers

  • Eric Christy showed some slides
    • The SHMS chambers performed well
    • There were 2 dead wires although it is though that these can be recovered
    • t0 offsets per wire have been implemented and the residuals have improved as expected
    • The two planes with low efficiencies contained the missing wires
    • Drift time along the wire is not completely understood at the moment, this will be looked into at a later date
    • The residuals have been reduced to 300 um
    • Ideally after alignment they expect to obtain a resolution on the order of 150 um

Hodoscopes

  • Simona will be trained by Sam on how to mount the PMT's on the quartz bars so that a ~200 micron (or thinner) cookie is obtained - this should happen during the second half of May
    • It was suggested that Simona should ask Sam to check on the expiration date of the glue
  • Simona will measure the quantum efficiency of the new UV glass PMTs; Carl Zorn will help Simona setting up a test stand for the QE measurements
  • She is working with a vendor to commit to a particular spec for the UV glass, however their typical a unicorn
  • Steve finished setting up the DAQ for the cosmic test of the quartz bars; Simona is working on setting up the trigger
  • Lead time on the new tubes is 12 weeks

Heavy Gas Cherenkov

  • No update provided

Aerogel Cerenkov

  • No update provided

Shower/Pre-shower

  • Dave Mack showed some slides regarding the HERMES XP3461 base modeling, see his slides
    • This is a follow up of the slides he showed last DWG meeting
    • He plans to model the gain and compare the difference between. the XP3462s
  • Vardan showed a slide from his logbook entry
    • SHMS calibration was C++ code written based on HMS and SOS Fortran code
    • The resolution improves for run 484 where the HV was increased by 100V
    • The SHMS calorimeter is behaving as expected
  • Hamlet commented that the resolutions we are obtaining are much better than what was obtained by the HERMES collaboration